Steady progress has been made towards the completion of Step IV. Examples of the tasks completed this week include: load tests on the PRY tie-downs; installation of the new optical fibres for the target and the junction boxes inside and outside the sync; and the fitting of cable-tray covers to the South Mezzanine. Work in Rack Room 2 progresses steadily too with the dump-circuits wired in and many of the large cables terminated.
The PPS annual functional test was carried out on Wednesday. This was a long day for those involved (Craig Macwaters, Henry Nebrensky and Phil Warburton). The test went well and the system was passed as functionally correct. A number of issues have been identified which will need to be addressed in the coming weeks.
Work to define, and then address, the issues noted in the 21Jan15 mock data-run is in hand. Chris Rogers completed the MICE Note summarising the run and documenting the issues to be addressed (MICE Note 456). Pierrick Hanlet as MOM is now helping to organise the work required to address the issues; particularly those for which coordination with the construction team or between detector groups is required. Its encouraging to note that a number of issues have already been resolved and that the number of issues that require cross-team coordination is low.
The near future:¶
The first set of PRY legs are now expected to arrive at RAL next week. The receipt and installation of the legs on the north side will then begin. Once the north-side installation has been completed, the legs will be removed so that the magnets can be moved to the parked position to allow the legs to be mounted in their final Step IV positions on the south side.
The PRY plates are scheduled to arrive at RAL with the second set of legs on 16Mar15. To expedite the installation of the PRY, the project team have decided to begin two-shift working from 16Mar15. The necessary technical and engineering effort has been secured. Roy Preece, Andy Nichols, Jason Tarrant and John Govans are now making the detailed plans necessary to make this, the final push in the installation of Step IV, a success.
The knock on effect to the run proposed for the 08Mar15 is that it has been necessary to scale back the ambitions. The plans for operations until the end of User Run 2015(01) are now:
○ MICE has been allocated two shifts (08:30-16:30 and 16:30-24:00) during which we may take beam into the Hall. ○ In view of the construction schedule and pressure on individuals it has been agreed to take data using the beam-line instrumentation as far as TOF I. • The positions of the instrumentation, the magnets and other equipment downstream of TOF I will be as required for the installation of the PRY. TOF1, and the equipment upstream of TOF1 is unaffected by the PRY activities. ○ The revised goals for operation on the 08Mar15 are to: • Time-in and commission the MICE target in the 700 MeV ISIS beam; • Time-in and commission the MICE DAQ, controls etc. and detectors up to TOF1; • Leave the experiment in such a state that alignment/calibration of the detectors up to TOF1 could begin. ○ Planning for the day so far is: • Lock and search Hall at the end of the construction work on Friday 06Mar15; • 08:00 on 08Mar16: □ Prepare for operation; • 10:00—22:00 on 08Mar16: □ Operation of the experiment; the run plan will be prepared by the MOM; • 22:00 (latest) on 08Mar16: □ Operations cease to allow … • 22:00-23:59 on 08Mar16: □ Close down Hall ready so that it is ready for the construction team to start on Monday 09Mar15.